What is the pH at 25˚C of a 1.000 L aqueous solution containing 1.90 moles of HBrO and 0.200 moles of LiBrO? Assume that LiBrO is a strong electrolyte.
Answer #1


Acid dissociation constant of HBrO, K, = 2.8x10-9 pKa = – logKa = - log (2.8x109) = 9.0– log 2.8 = 9.0-0.45 = 8.55 According